OpenStudy (poofypenguin):

Given vectors a, b, x, and y: If 3x-2y=a and 5x-3y=b, express vectors x and y in terms of a and b.

OpenStudy (anonymous):

Multiply First equation by 3 and Second Equation by 2, 9x - 6y = 3a (First Equation) 10x - 6y = 2b (Second Equation) Subtract the two equations, -x = 3a - 2b Or, x = 2b - 3a.. Like wise, from first equation, 2y = 3x - a So, 2y = 3(2b - 3a) - a 2y = 6b - 9a - a y = (6b - 10a)/2 So, y = 3b - 5a... Therefore, x = (2b - 3a) and y = (3b - 5a) is the answer...

OpenStudy (poofypenguin):

Thanks for the solution!:D

OpenStudy (anonymous):

Welcome dear...